a person is performing a pulmonary function test and is asked to inhale with maximum effort. His lung volume increases from 2.4 to 6.2 L. If his lungs started out at 1.0 atm pressure, what would the pressure be if air molecules did not rush in upon inhalation?



1
Expert's answer
2021-04-13T02:08:02-0400

P1V1 = P2V2

= 2.4 × 1 = 6.2 × X

= 2.4 = 6.2X

= 0.387atm